# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 6. Bounding the tree-child SNPR distance
First suppose that ${\mathcal{N}}^{\prime\prime}$ contains a pair of parallel edges. Then $(p_{u},c_{u})$, $(p_{u},u)$, and $(u,c_{u})$ are edges of an underlying three-cycle in ${\mathcal{N}}$. Assumption (A) and Lemma 4.1(i) imply that all three edges incident with $c_{u}$ are edges in $R_{u}$. If $(p_{u},u)\notin R_{u}$, then $u$ has in-degree zero and out-degree one in $R_{u}$. It follows that $u$ is a vertex of ${\mathcal{R}}$ but not a vertex of the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}$. Hence, the unique child of $u$ in $R_{u}$ has in-degree one in $R_{u}$ because ${\mathcal{R}}$ is a root extension of ${\mathcal{D}}$, a contradiction as $c_{u}$ has in-degree two in $R_{u}$. Thus, $(p_{u},u)\in R_{u}$. It follow that $D_{u}$ contains a pair of parallel edges because $e\notin{\mathcal{R}}$, a contradiction to ${\mathcal{D}}$ being tree-child.
@@
Second suppose that ${\mathcal{N}}^{\prime\prime}$ contains an edge that is incident with two reticulations. Then $p_{u}$ and $c_{u}$ are reticulations in ${\mathcal{N}}$. It follows from (A) and Lemma 4.1(i), that $R_{u}$ contains the three edges incident with $c_{u}$ and the three edges incident with $p_{u}$. Thus, $D_{u}$ contains an edge that is incident with two reticulations because $e\notin{\mathcal{R}}$, another contradiction.
@@
Third suppose that ${\mathcal{N}}^{\prime\prime}$ contains a pair of sibling reticulations. Then $c_{u}$ is a reticulation and $p_{u}$ is a tree vertex whose child that is not $u$, say $s_{u}$, is a reticulation in ${\mathcal{N}}$. Again by (A) and Lemma 4.1(i), $R_{u}$ contains all three edges that are incident with $c_{u}$ and there exists an element $R_{u^{\prime}}\in{\mathcal{R}}$ with $u^{\prime}\in\{\rho,1,2,\ldots,k\}$ such that $R_{u^{\prime}}$ contains all three edges incident with $s_{u}$. If $u\neq u^{\prime}$, then $(p_{u},u)\notin{\mathcal{R}}$ and, thus, $u$ has in-degree zero and out-degree one in $R_{u}$. It follows that $u$ is a vertex of ${\mathcal{R}}$ but not a vertex of the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}$. Hence, the unique child of $u$ in $R_{u}$ has in-degree one in $R_{u}$, a contradiction as $c_{u}$ has in-degree two in $R_{u}$. We may therefore assume that $u=u^{\prime}$. But then $R_{u}$ contains a pair of sibling reticulations $s_{u}$ and $c_{u}$ because $e\notin{\mathcal{R}}$, a final contradiction. ∎
@@
It now follows from (6.4.4) and (6.4.5) and the construction as detailed in (C1)–(C3) that ${\mathcal{N}}^{\prime\prime}$ is a tree-child network on $X$ that can be obtained from ${\mathcal{N}}$ by a single SNPR±. We next show that ${\mathcal{D}}^{\prime}$ is a phylogenetic digraph for ${\mathcal{N}}^{\prime\prime}$. To this end, we construct a root extension of ${\mathcal{D}}^{\prime}$ in ${\mathcal{N}}^{\prime\prime}$.
@@
If ${\mathcal{N}}^{\prime\prime}$ has been obtained from ${\mathcal{N}}$ as described in (C1), obtain a root extension $R_{ij}$ of $D_{ij}^{\prime}$ from $R_{i}$ and $R_{j}$ by subdividing $f$ in $R_{j}$ with a new vertex $p$ and adding the edge $(p,w)$. Otherwise, if ${\mathcal{N}}^{\prime\prime}$ has been obtained from ${\mathcal{N}}$ as described in (C2) or (C3), obtain $R_{ij}$ from $R_{i}$ and $R_{j}$ by adding the edge $(p,v)$, where $p$ is a new vertex, and adding the edges $(p,v)$ and $(p,w)$. Then, as ${\mathcal{R}}$ is a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ and $u$ is a tree vertex in ${\mathcal{N}}$ by (6.4.1), it follows that the digraph obtained from
′′=(${\mathcal{R}}^{\prime\prime}=({\mathcal{R}}-\{R_{i},R_{j}\})\cup\{R_{ij}\}$
@@
By combining (6.4.3) and (6.4.6), it now follows from the induction assumption that
$d_{\rm tc}({\mathcal{N}}^{\prime\prime},{\mathcal{N}}^{\prime})\leq c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1.$
Hence, there exists a tree-child SNPR sequence $\sigma$ connecting ${\mathcal{N}}^{\prime\prime}$ and ${\mathcal{N}}^{\prime}$ with $w(\sigma)\leq c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1$. Since ${\mathcal{N}}^{\prime\prime}$ can be obtained from ${\mathcal{N}}$ by a single SNPR±, we have
$d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leq 2+c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1=c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}.$
@@
Figure 3. Example of two tree-child networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ and an agreement tree-child digraph ${\mathcal{D}}$ for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ with $86=d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})<m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=$ 8. An extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ is indicated by thick lines.
Figure 3 shows two tree-child networks for which the inequality established in Lemma 6.4 is strict. However, the next lemma shows that, for two tree-child networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, the difference $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})-d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ cannot be arbitrary large. In preparation for the lemma, we need an additional definition. Let ${\mathcal{D}}$ be a phylogenetic digraph of a phylogenetic network ${\mathcal{N}}$ on $X$. Furthermore, let ${\mathcal{R}}$ be an extension of ${\mathcal{D}}$ in ${\mathcal{N}}$, and let ${\mathcal{M}}$ be the embedding that underlies ${\mathcal{R}}$. Now consider a directed path $P$ in ${\mathcal{M}}$. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ be the subset of reticulations in ${\mathcal{N}}$ that lie on $P$. Then the path extension of $P$ contains precisely all edges of $P$ and, additionally, each edge of a maximal length directed path in ${\mathcal{N}}$ that only consists of edges in $E_{\mathcal{R}}-E_{\mathcal{M}}$ and ends at a vertex in $V$. Note that the path extension of $P$ may contain each edge of $P$ and no additional edge, even if $V\neq\emptyset$.
Lemma 6.5.
@@
Let $\sigma=({\mathcal{N}}={\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{t}={\mathcal{N}}^{\prime})$ be a tree-child SNPR sequence connecting ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ such that ${\mathcal{N}}_{i-1}$ and ${\mathcal{N}}_{i}$ are non-isomorphic for each $i\in\{1,2,\ldots,t\}$. It follows from [6, Proposition 3.2] that $\sigma$ exists. By Corollary 6.3, we may assume that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by an SNPR+ or an SNPR±, or ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- for each $i\in\{1,2,\ldots,t\}$. If the latter holds, then, by the reversibility of SNPR, $({\mathcal{N}}^{\prime}={\mathcal{N}}_{t},\ldots,{\mathcal{N}}_{2},{\mathcal{N}}_{1},{\mathcal{N}}_{0}={\mathcal{N}})$ is a tree-child SNPR sequence connecting ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ and ${\mathcal{N}}_{i}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR+ for each $i\in\{t-1,t-2,\ldots,0\}$. Hence, we may assume without loss of generality that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by either an SNPR+ or an SNPR±. We show by induction on $t$ that there exist an agreement tree-child digraph ${\mathcal{D}}$ for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ and extensions ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively, such that $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}’)\leq w(\sigma)$. The lemma then follows by choosing $\sigma$ such that $w(\sigma)=d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$.
@@
If $t=1$, then there are two cases to consider. First assume that ${\mathcal{N}}^{\prime}$ can be obtained from ${\mathcal{N}}$ by an SNPR+. Then $w(\sigma)=1$, and ${\mathcal{N}}$ is an agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. Trivially, there is an extension ${\mathcal{R}}$ of ${\mathcal{N}}$ in ${\mathcal{N}}$ and an extension ${\mathcal{R}}^{\prime}$ of ${\mathcal{N}}$ in ${\mathcal{N}}^{\prime}$ such that $|E_{\mathcal{N}}-E_{\mathcal{R}}|+|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|=1$ and, thus, $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leq\frac{1}{2}\cdot 1<w(\sigma)$. Second assume that ${\mathcal{N}}^{\prime}$ can be obtained from ${\mathcal{N}}$ by an SNPR± in which case $w(\sigma)=2$. Recalling the global assumption stated at the end of Section 5, let $e=(u,v)$ be the tree edge in ${\mathcal{N}}$ that is deleted in the process of obtaining ${\mathcal{N}}^{\prime}$ from ${\mathcal{N}}$. Let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ in ${\mathcal{N}}$ that is not $v$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. If $p_{u}$ is a tree vertex, let $s$ be the child of $p_{u}$ in ${\mathcal{N}}$ that is not $u$. Furthermore, if $c_{u}$ is a reticulation, let $s^{\prime}$ be the parent of $c_{u}$ in ${\mathcal{N}}$ that is not $u$. If $s$ and $c_{u}$ are both reticulations, let ${\mathcal{D}}$ be the leaf-labelled acyclic digraph ${\mathcal{D}}$ obtained from ${\mathcal{N}}$ by deleting $e$ and $(s^{\prime},c_{u})$, and suppressing $u$, $c_{u}$, and $s^{\prime}$. Otherwise, if at least one of $s$ and $c_{u}$ is not a reticulation, let ${\mathcal{D}}$ be the leaf-labelled acyclic digraph ${\mathcal{D}}$ obtained from ${\mathcal{N}}$ by deleting $e$ and suppressing $u$. In both cases, ${\mathcal{D}}$ is an agreement digraph of ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. We next show that ${\mathcal{D}}$ is tree-child. If ${\mathcal{D}}$ contains a pair of parallel edges or a stack, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$, a contradiction to the choice of $\sigma$. On the other hand, if ${\mathcal{D}}$ contains a pair of sibling reticulations, then $s$ and $c_{u}$ are reticulations in ${\mathcal{N}}$. By construction, it follows that there is no reticulation in ${\mathcal{D}}$ that corresponds to $c_{u}$. Hence, ${\mathcal{D}}$ is tree-child. Moreover, there are extensions ${\mathcal{R}}$ of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{R}}^{\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ such that $|E_{\mathcal{N}}-E_{\mathcal{R}}|+|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|\leq 2+2=4$, where the first inequality becomes an equality only if $s$ and $c_{u}$ are both reticulations in ${\mathcal{N}}$. It now follows that $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leq\frac{1}{2}\cdot 4=w(\sigma)$. This completes the proof of the base case.
@@
By Corollary 6.3, we may again assume that ${\mathcal{N}}_{t-1}$ can be obtained from ${\mathcal{N}}_{t-2}$ by an SNPR+ or an SNPR±, or ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- for each $i\in\{1,2,\ldots,t-1\}$. Observe that $w(\sigma)=w(\sigma_{1})+w(\sigma_{2})$. By the induction assumption, we have
$\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{0},{\mathcal{N}}_{t-1})\leq w(\sigma_{1}).$
@@
(1) $\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{0},{\mathcal{N}}_{t-1})=\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|)\leq w(\sigma_{1}).$
@@
Assume that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by an SNPR+, in which case $w(\sigma_{2})=1$. Let $(u,w)$ and $(u^{\prime},w^{\prime})$ be the two edges in ${\mathcal{N}}_{t-1}$ that are subdivided with a new vertex $v$ and $v^{\prime}$, respectively, in obtaining ${\mathcal{N}}_{t}$. Since ${\mathcal{N}}_{t}$ is tree-child, $(u,w)$ and $(u^{\prime},w^{\prime})$ are tree edges. Furthermore, either $(v,v^{\prime})$ or $(v^{\prime},v)$ is a reticulation edge in ${\mathcal{N}}_{t}$. Without loss of generality, we may assume that $(v^{\prime},v)$ is a reticulation edge in ${\mathcal{N}}_{t}$. It now follows from Lemma 6.1(ii) that ${\mathcal{D}}^{\prime}$ is also a tree-child digraph for ${\mathcal{N}}_{t}$ and there exists an extension ${\mathcal{R}}_{t}^{\prime}$ of ${\mathcal{D}}^{\prime}$ in ${\mathcal{N}}_{t}$ such that
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$
@@
First assume that $e\notin{\mathcal{R}}^{\prime}_{t-1}$. It follows from Lemma 6.1(iii) that ${\mathcal{D}}^{\prime}$ is also a tree-child digraph for ${\mathcal{N}}_{t}$ and there exists an extension ${\mathcal{R}}^{\prime}_{t}$ of ${\mathcal{D}}^{\prime}$ in ${\mathcal{N}}_{t}$ such that
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$$\leq$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$

### v2

@@ 6 Bounding the tree-child SNPR distance
First suppose that $\mathcal{N}$$^{\prime\prime}$ contains a pair of parallel edges. Then $(p_{u},c_{u})$, $(p_{u},u)$, and $(u,c_{u})$ are edges of an underlying three-cycle in $\mathcal{N}$}. Assumption (A) and Lemma 5(i) imply that all three edges incident with $c_{u}$ are edges in $R_{u}$. If $(p_{u},u)\notin R_{u}$, then $u$ has in-degree zero and out-degree one in $R_{u}$. It follows that $u$ is a vertex of $\mathcal{R}$} but not a vertex of the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$}. Hence, the unique child of $u$ in $R_{u}$ has in-degree one in $R_{u}$ because $\mathcal{R}$} is a root extension of $\mathcal{D}$}, a contradiction as $c_{u}$ has in-degree two in $R_{u}$. Thus, $(p_{u},u)\in R_{u}$. It follow that $D_{u}$ contains a pair of parallel edges because $e\notin{\mathcal{R}}$, a contradiction to $\mathcal{D}$} being tree-child.
@@
Second suppose that $\mathcal{N}$$^{\prime\prime}$ contains an edge that is incident with two reticulations. Then $p_{u}$ and $c_{u}$ are reticulations in $\mathcal{N}$}. It follows from (A) and Lemma 5(i), that $R_{u}$ contains the three edges incident with $c_{u}$ and the three edges incident with $p_{u}$. Thus, $D_{u}$ contains an edge that is incident with two reticulations because $e\notin{\mathcal{R}}$, another contradiction.
@@
Third suppose that $\mathcal{N}$$^{\prime\prime}$ contains a pair of sibling reticulations. Then $c_{u}$ is a reticulation and $p_{u}$ is a tree vertex whose child that is not $u$, say $s_{u}$, is a reticulation in $\mathcal{N}$}. Again by (A) and Lemma 5(i), $R_{u}$ contains all three edges that are incident with $c_{u}$ and there exists an element $R_{u^{\prime}}\in{\mathcal{R}}$ with $u^{\prime}\in\{\rho,1,2,\ldots,k\}$ such that $R_{u^{\prime}}$ contains all three edges incident with $s_{u}$. If $u\neq u^{\prime}$, then $(p_{u},u)\notin{\mathcal{R}}$ and, thus, $u$ has in-degree zero and out-degree one in $R_{u}$. It follows that $u$ is a vertex of $\mathcal{R}$} but not a vertex of the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$}. Hence, the unique child of $u$ in $R_{u}$ has in-degree one in $R_{u}$, a contradiction as $c_{u}$ has in-degree two in $R_{u}$. We may therefore assume that $u=u^{\prime}$. But then $R_{u}$ contains a pair of sibling reticulations $s_{u}$ and $c_{u}$ because $e\notin{\mathcal{R}}$, a final contradiction. ∎
@@
It now follows from (13.4) and (13.5) and the construction as detailed in (C1)–(C3) that $\mathcal{N}$$^{\prime\prime}$ is a tree-child network on $X$ that can be obtained from $\mathcal{N}$} by a single SNPR±. We next show that $\mathcal{D}$$^{\prime}$ is a phylogenetic digraph for $\mathcal{N}$$^{\prime\prime}$. To this end, we construct a root extension of $\mathcal{D}$$^{\prime}$ in $\mathcal{N}$$^{\prime\prime}$.
@@
If $\mathcal{N}$$^{\prime\prime}$ has been obtained from $\mathcal{N}$} as described in (C1), obtain a root extension $R_{ij}$ of $D_{ij}^{\prime}$ from $R_{i}$ and $R_{j}$ by subdividing $f$ in $R_{j}$ with a new vertex $p$ and adding the edge $(p,w)$. Otherwise, if $\mathcal{N}$$^{\prime\prime}$ has been obtained from $\mathcal{N}$} as described in (C2) or (C3), obtain $R_{ij}$ from $R_{i}$ and $R_{j}$ by adding the edge $(p,v)$, where $p$ is a new vertex, and adding the edges $(p,v)$ and $(p,w)$. Then, as $\mathcal{R}$} is a root extension of $\mathcal{D}$} in $\mathcal{N}$} and $u$ is a tree vertex in $\mathcal{N}$} by (13.1), it follows that the digraph obtained from
${\mathcal{R}}^{\prime\prime}=({\mathcal{R}}-\{R_{i},R_{j}\})\cup\{R_{ij}\}$
@@
By combining (13.3) and (13.6), it now follows from the induction assumption that
$d_{\rm tc}({\mathcal{N}}^{\prime\prime},{\mathcal{N}}^{\prime})\leqslant c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1.$
Hence, there exists a tree-child SNPR sequence $\sigma$ connecting $\mathcal{N}$$^{\prime\prime}$ and $\mathcal{N}$$^{\prime}$ with $w(\sigma)\leqslant c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1$. Since $\mathcal{N}$$^{\prime\prime}$ can be obtained from $\mathcal{N}$} by a single SNPR±, we have
$d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leqslant 2+c_{\mathcal{D}}-1+c_{\mathcal{D}}^{\prime}-1=c_{\mathcal{D}}+c_{\mathcal{D}}^{\prime}.$
@@
Figure 9: An example of two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ and an agreement tree-child digraph $\mathcal{D}$} for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ for which $6=d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ but $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})=8$. An extension of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ is indicated by red lines.
Figure 9 shows two tree-child networks for which the inequality established in Lemma 13 is strict. However, the next lemma shows that, for two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, the difference $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})-d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ cannot be arbitrary large. In preparation for the lemma, we need an additional definition. Let $\mathcal{D}$} be a phylogenetic digraph of a phylogenetic network $\mathcal{N}$} on $X$. Furthermore, let $\mathcal{R}$} be an extension of $\mathcal{D}$} in $\mathcal{N}$}, and let $\mathcal{M}$} be the embedding that underlies $\mathcal{R}$}. Now consider a directed path $P$ in $\mathcal{M}$}. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ be the subset of reticulations in $\mathcal{N}$} that lie on $P$. Then the path extension of $P$ contains precisely all edges of $P$ and, additionally, each edge of a maximal length directed path in $\mathcal{N}$} that only consists of edges in $E_{\mathcal{R}}-E_{\mathcal{M}}$ and ends at a vertex in $V$. Note that the path extension of $P$ may contain each edge of $P$ and no additional edge, even if $V\neq\emptyset$.
Lemma 14.
@@
Let $\sigma=({\mathcal{N}}={\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{t}={\mathcal{N}}^{\prime})$ be a tree-child SNPR sequence connecting $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ such that ${\mathcal{N}}_{i-1}$ and ${\mathcal{N}}_{i}$ are non-isomorphic for each $i\in\{1,2,\ldots,t\}$. It follows from Bordewich et al. [7, Proposition 3.2] that $\sigma$ exists. By Corollary 12, we may assume that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by an SNPR+ or an SNPR±, or ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- for each $i\in\{1,2,\ldots,t\}$. If the latter holds, then, by the reversibility of SNPR,
$({\mathcal{N}}^{\prime}={\mathcal{N}}_{t},\ldots,{\mathcal{N}}_{2},{\mathcal{N}}_{1},{\mathcal{N}}_{0}={\mathcal{N}})$
is a tree-child SNPR sequence connecting $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ and ${\mathcal{N}}_{i}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR+ for each $i\in\{t-1,t-2,\ldots,0\}$. Hence, we may assume without loss of generality that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by either an SNPR+ or an SNPR±. We show by induction on $t$ that there exist an agreement tree-child digraph $\mathcal{D}$} for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ and extensions $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$ of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, respectively, such that $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}’)\leqslant w(\sigma)$. The lemma then follows by choosing $\sigma$ such that $w(\sigma)=d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$.
@@
If $t=1$, then there are two cases to consider. First assume that $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by an SNPR+. Then $w(\sigma)=1$, and $\mathcal{N}$} is an agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. Trivially, there is an extension $\mathcal{R}$} of $\mathcal{N}$} in $\mathcal{N}$} and an extension $\mathcal{R}$$^{\prime}$ of $\mathcal{N}$} in $\mathcal{N}$$^{\prime}$ such that $|E_{\mathcal{N}}-E_{\mathcal{R}}|+|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|=1$ and, thus, $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leqslant\frac{1}{2}\cdot 1<w(\sigma)$. Second assume that $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by an SNPR± in which case $w(\sigma)=2$. Recalling the global assumption stated at the end of Section 5, let $e=(u,v)$ be the tree edge in $\mathcal{N}$} that is deleted in the process of obtaining $\mathcal{N}$$^{\prime}$ from $\mathcal{N}$}. Let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ in $\mathcal{N}$} that is not $v$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. If $p_{u}$ is a tree vertex, let $s$ be the child of $p_{u}$ in $\mathcal{N}$} that is not $u$. Furthermore, if $c_{u}$ is a reticulation, let $s^{\prime}$ be the parent of $c_{u}$ in $\mathcal{N}$} that is not $u$. If $s$ and $c_{u}$ are both reticulations, let $\mathcal{D}$} be the leaf-labelled acyclic digraph $\mathcal{D}$} obtained from $\mathcal{N}$} by deleting $e$ and $(s^{\prime},c_{u})$, and suppressing $u$, $c_{u}$, and $s^{\prime}$. Otherwise, if at least one of $s$ and $c_{u}$ is not a reticulation, let $\mathcal{D}$} be the leaf-labelled acyclic digraph $\mathcal{D}$} obtained from $\mathcal{N}$} by deleting $e$ and suppressing $u$. In both cases, $\mathcal{D}$} is an agreement digraph of $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. We next show that $\mathcal{D}$} is tree-child. If $\mathcal{D}$} contains a pair of parallel edges or a stack, then ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$, a contradiction to the choice of $\sigma$. On the other hand, if $\mathcal{D}$} contains a pair of sibling reticulations, then $s$ and $c_{u}$ are reticulations in $\mathcal{N}$}. By construction, it follows that there is no reticulation in $\mathcal{D}$} that corresponds to $c_{u}$. Hence, $\mathcal{D}$} is tree-child. Moreover, there are extensions $\mathcal{R}$} of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{R}$$^{\prime}$ of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ such that $|E_{\mathcal{N}}-E_{\mathcal{R}}|+|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|\leqslant 2+2=4$, where the first inequality becomes an equality only if $s$ and $c_{u}$ are both reticulations in $\mathcal{N}$}. It now follows that $\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})\leqslant\frac{1}{2}\cdot 4=w(\sigma)$. This completes the proof of the base case.
@@
By Corollary 12, we may again assume that ${\mathcal{N}}_{t-1}$ can be obtained from ${\mathcal{N}}_{t-2}$ by an SNPR+ or an SNPR±, or ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- for each $i\in\{1,2,\ldots,t-1\}$. Observe that $w(\sigma)=w(\sigma_{1})+w(\sigma_{2})$. By the induction assumption, we have
$\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{0},{\mathcal{N}}_{t-1})\leqslant w(\sigma_{1}).$
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}}_{0},{\mathcal{N}}_{t-1})=\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t-1}}-E_{{\mathcal{R}}^{\prime}_{t-1}}|)\leqslant w(\sigma_{1}).$ (1)
@@
Assume that ${\mathcal{N}}_{t}$ can be obtained from ${\mathcal{N}}_{t-1}$ by an SNPR+, in which case $w(\sigma_{2})=1$. Let $(u,w)$ and $(u^{\prime},w^{\prime})$ be the two edges in ${\mathcal{N}}_{t-1}$ that are subdivided with a new vertex $v$ and $v^{\prime}$, respectively, in obtaining ${\mathcal{N}}_{t}$. Since ${\mathcal{N}}_{t}$ is tree-child, $(u,w)$ and $(u^{\prime},w^{\prime})$ are tree edges. Furthermore, either $(v,v^{\prime})$ or $(v^{\prime},v)$ is a reticulation edge in ${\mathcal{N}}_{t}$. Without loss of generality, we may assume that $(v^{\prime},v)$ is a reticulation edge in ${\mathcal{N}}_{t}$. It now follows from Lemma 10(ii) that $\mathcal{D}$$^{\prime}$ is also a tree-child digraph for ${\mathcal{N}}_{t}$ and there exists an extension ${\mathcal{R}}_{t}^{\prime}$ of $\mathcal{D}$$^{\prime}$ in ${\mathcal{N}}_{t}$ such that
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ $\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$
@@
First assume that $e\notin{\mathcal{R}}^{\prime}_{t-1}$. It follows from Lemma 10(iii) that $\mathcal{D}$$^{\prime}$ is also a tree-child digraph for ${\mathcal{N}}_{t}$ and there exists an extension ${\mathcal{R}}^{\prime}_{t}$ of $\mathcal{D}$$^{\prime}$ in ${\mathcal{N}}_{t}$ such that
@@
$\frac{1}{2}m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ $\leqslant$ $\frac{1}{2}(|E_{{\mathcal{N}}_{0}}-E_{{\mathcal{R}}^{\prime}_{0}}|+|E_{{\mathcal{N}}_{t}}-E_{{\mathcal{R}}^{\prime}_{t}}|)$